OpenStudy (meilendurcer):

Solve for x: 2|x - 3| + 1 = 7 A. x = 0, x = 6 B. x = -1, x = 6 C. x = 0, x = 9 D. No solutions I think it's either B or D

OpenStudy (anonymous):

Option A is correct...

OpenStudy (anonymous):

D no solutions

OpenStudy (anonymous):

9,6,0 doesn't work neither does -1

OpenStudy (anonymous):

In the place of x substitute 0 or 6 it satisfies the given eqution...

OpenStudy (anonymous):

your right.

OpenStudy (meilendurcer):

ahh i see k thanks!

OpenStudy (anonymous):

\[2\left| x-3 \right|+1=7,2\left| x-3 \right|=7-1=6,\left| x-3 \right|=3\] \[x-3=\pm 3,x=3+3=6,x=-3+3=0\]